OpenStudy (mony01):

If 27 J of work are needed to stretch a spring from 16 cm to 22 cm and 45 J are needed to stretch it from 22 cm to 28 cm, what is the natural length of the spring?

OpenStudy (dan815):

do u know the equation for energy stored in a spring

OpenStudy (dan815):

F=kx

OpenStudy (anonymous):

E = (1/2) K * x^2, where K is the spring constant and x is the displacement

OpenStudy (dan815):

integrate to find work so w=int(kx dx)

OpenStudy (dan815):

plug in your values and solve

OpenStudy (mony01):

do i subtract 22 from 16 first?

OpenStudy (dan815):

http://www.sosmath.com/CBB/viewtopic.php?t=987

OpenStudy (dan815):

here u go a full solution http://www.freemathhelp.com/forum/threads/46261-Natural-length-of-a-spring-given-work

OpenStudy (dan815):

we dont know how much it was stretched past the equlibrium position

OpenStudy (dan815):

w=1/2k*(x)^2 where x is the stretch from equilibirum

OpenStudy (dan815):

w=(k/2)*(22-a)^2-[k/2*(16-a)^2]=27 w=(k/2)*(28-a)^2-[k/2*(22-a)^2]=45

OpenStudy (dan815):

you have 2 equations 2 unknows solve them

OpenStudy (mony01):

would the answer be 9?

OpenStudy (whpalmer4):

does 9 give a valid solution if you plug it back in to both equations?

OpenStudy (mony01):

where can i plug in 9 at a?

OpenStudy (whpalmer4):

well, how did you come up with 9? surely it was on opposite side of an equals sign from a letter, right? which letter? :-)

OpenStudy (whpalmer4):

\[\frac{k}{2}(22-a)^2-\frac{k}{2}(16-a)^2=27\]\[\frac{k}2(28-a)^2-\frac{k}2(22-a)^2=45\]

OpenStudy (dan815):

have learnt how to solve equations

OpenStudy (dan815):

think of it like balancing wieghts

OpenStudy (whpalmer4):

\[(16-a)^2 = a^2-32a+256\]\[(22-a)^2=a^2-44a+484\]\[(28-a)^2=a^2-56a+784\] Multiply both equations by \((2/k)\): \[(22-a)^2-(16-a)^2 = 54/k\]\[(28-a)^2-(22-a)^2 = 90/k\]\[a^2-44a+484-a^2+32a-256 = 54/k\]\[a^2-56a+784-a^2+44a-484=90/k\] \[-12a+228=54/k\]\[-12a+300=90/k\]Multiply one row by -1 and add together \[72=36/k\]You should be able to find \(a\)...